In a historical movie, two knights on horseback start from rest 89 m apart and ride directly toward each other to do battle. Sir George's acceleration has a magnitude of 0.17 m/s2, while Sir Alfred's has a magnitude of 0.26 m/s2. Relative to Sir George's starting point, where do the knights collide
Answer #1

SOLUTION :


Let they collide after t secs. 


So,


Distance travelled by Sir George = 1/2 a t^2 = 1/2 * 0.17 * t^2 = 0.085 t^2


Distance travelled by Sir Alfred = 1/2 a’ t^2 = 1/2 * 0.26* t^2 = 0.13 t^2


As they collide, the sum of their distances travelled = 89 m

=> 0.085 t^2 + 0.13 t^2 = 89

=> 0.215 t^2 = 89

=> t^2 = 89 / 0.215 = 413.9535 


Sir George ’s distance travelled = 0.085 * 413.9535 = 35.19 m approx. 


Hence,  they collide at a  point 35.19  m  approx. from the starting point of Sir George. (ANSWER).

answered by: Tulsiram Garg